# Reproducing the first and second moment of empirical degree distributions

### v3

@@ VI.3 Reproducing a network structure
A compact indicator of a model performance in reconstructing a network structure is provided by the Bayesian Information Criterion (BIC), reading
@@
where V=N(N1)/2V=N(N-1)/2 accounts for the system size, the number of parameters amounts at $c=N$ for the UBCM, $c=1$ for the dcGM, $c=2$ for the fit2SM and the log-likelihood reads
@@
$\delta\text{BIC}_{model}=\frac{\text{BIC}_{model}-\text{BIC}_{min}}{\text{BIC}_{model}},$ (27)
where $\text{BIC}_{min}=\min_{model}\{\text{BIC}_{model}\}$; while a null value of such an index identifies the best-performing model on the considered snapshot, the larger its value, the worse the performance of the corresponding model. As fig. 4 shows, our canonical, non-linear model
  • steadily outperforms both the dcGM and the (much) more demanding UBCM at the daily and weekly time-scales, i.e. on sparser configurations (in fact, $\text{BIC}_{\text{fit2SM}}<\text{BIC}_{\text{dcGM}}$ on $90.76\%$ of daily snapshots and $97.63\%$ of weekly snapshots; $\text{BIC}_{\text{fit2SM}}<\text{BIC}_{\text{UBCM}}$ on $99.97\%$ of daily snapshots and $100.00\%$ of weekly snapshots);
@@
Overall, these findings suggest the presence of a dependency between the performance of the fit2SM and the network link density, i.e. the sparser the network, the better its performance. To test such an hypothesis, let us scatter the index of relative performance
$\Delta\text{BIC}_{model}=\frac{\text{BIC}_{model}-\text{BIC}_{\text{fit2SM}}}{\text{BIC}_{model}}$ (28)
@@
$\Delta S_{model}=\frac{\langle S\rangle_{model}-\langle S\rangle_{\text{fit2SM}}}{\langle S\rangle_{model}}=\frac{\langle S\rangle_{model}-S^{*}}{\langle S\rangle_{model}}$ (29)
for both the UBCM and the dcGM. As fig. 5 reveals, whenever $\Delta S_{model}$ is large - typically the case for sparser configurations - model selection favours the fit2SM.
@@ VII Discussion
After reviewing a number of negative results about the possibility of defining an ERG constraining both the degree sequence and the total number of two-stars, we propose a minimal, canonical model capable of reproducing the variance of a degree distribution while accounting for the nodes heterogeneity: in order to achieve such a goal, a fitness-induced variant of the 2SM, named fit2SM (i.e. a ‘softened’ non-linear ERG whose definition is inspired to its fitness-based, linear counterpart), has been employed.
Besides achieving a large accuracy in reproducing a network structure at different levels, the fit2SM seems also to mitigate a problem affecting the dcGM, i.e. that of returning sampled configurations with too many isolated nodes: in fact, $\langle N_{\text{fit2SM}}^{0}\rangle<\langle N_{\text{dcGM}}^{0}\rangle$ on $97.63\%$ of daily snapshots, $98.61\%$ of weekly snapshots and $67.88\%$ of monthly snapshots; interestingly, $\langle N_{\text{fit2SM}}^{0}\rangle<\langle N_{\text{UBCM}}^{0}\rangle$ on $58.79\%$ of monthly snapshots and $38.10\%$ of quarterly snapshots but $\langle N_{\text{fit2SM}}^{0}\rangle<\langle N_{\text{UBCM}}^{0}\rangle$ on $66.67\%$ of monthly snapshots and $55.56\%$ of quarterly snapshots since 2009 (see also Appendix F).
@@
Future research will explore non-linear models enforcing more complex patterns (e.g. the Strauss one, constraining the total number of triangles).
@@ X Acknowledgments
DG and TS acknowledge support from the projects ‘Reconstruction, Resilience and Recovery of Socio-Economic Networks’ RECON-NET EP_FAIR_005 - PE0000013 ‘FAIR’, Funded by the European Union Next Generation EU, PNRR Mission 4 Component 2, Investment 3.1; ‘RE-Net: Reconstructing economic networks: from physics to machine learning and back’ - 2022MTBB22, Funded by the European Union Next Generation EU, PNRR Mission 4 Component 2 Investment 1.1, CUP: D53D23002330006; ‘C2T - From Crises to Theory: towards a science of resilience and recovery for economic and financial systems’ - P2022E93B8, Funded by the European Union Next Generation EU, PNRR Mission 4 Component 2 Investment 1.1, CUP: D53D23019330001. This work is also supported by the European Union - NextGenerationEU and funded by the Italian Ministry of University and Research (MUR) - National Recovery and Resilience Plan (Piano Nazionale di Ripresa e Resilienza, PNRR), project ‘A Multiscale integrated approach to the study of the nervous system in health and disease’ (MNESYS, PE0000006, DN. 1553, 11/10/2022). We thank Giulio Cimini for fruitful and insightful discussions.
@@ XII Competing Interests
We employ transaction-level records from the Electronic Market for Interbank Deposits (eMID), a screen-based market for unsecured deposits [8, 38, 39], restricting the analysis to the segment that represents the vast majority of the activity on such a market, i.e. the overnight one. Our sample spans trading days from January $1999$ to September $2012$: each trading day $d$ defines a weighted, directed matrix $\mathbf{V}(d)$ whose entry $v_{ij}(d)$ equals the total notional amount lent by bank $i$ to bank $j$ on day $d$; naturally, $v_{ii}(d)=0$.
Throughout the paper, the labels ‘daily’, ‘weekly’, ‘monthly’, ‘quarterly’ and ‘yearly’ refer to calendar aggregations of daily records: more specifically, ‘weekly’ refers to ISO calendar weeks (Monday to Sunday), ‘monthly’ to calendar months, ‘quarterly’ to standard quarters (Q1: Jan-Mar; Q2: Apr-Jun; Q3: Jul-Sep; Q4: Oct-Dec) and ‘yearly’ to calendar years. Importantly, trading on eMID only occurs on business days, so that weekends and bank holidays do not contribute to observations: consequently, the number of trading days within a given calendar window is not constant across windows of the same kind; in practice, we aggregate over the set of trading days that are present in the raw records during a specific window. Let $\Delta_{t}$ denote the set of trading days belonging to the calendar window indexed by $t$ (week, month, quarter or year). We, thus, aggregate weights according to
@@
as a last observation, let us stress that, for each window, we restrict the set of nodes to the banks that are active within that window, i.e. that are involved in at least one transaction during $\Delta_{t}$: the number of nodes, thus, becomes a time-dependent quantity.
@@
with $a_{ii}^{\Delta}(t)=0$. Naturally, we compute the node strengths from the underlying, symmetrised weights as
@@
To ease the mathematical manipulations, let us focus on binary, undiretced networks (BUNs). Linear ERGs are described by Hamiltonians reading
@@
hence inducing models that can be factorised as
@@
hence inducing models that can be factorised as well as
@@
where $\mathbf{P}=\langle\mathbf{A}\rangle$. The mean-field approximation consists precisely in replacing $\mathbf{A}$ with $\mathbf{P}$, inside $f_{ij}$, in eq. 36.
@@
$\sum_{i}\binom{\langle k_{i}\rangle}{2}=\frac{1}{2}\sum_{i}[\langle k_{i}\rangle^{2}-\langle k_{i}\rangle]=\frac{1}{2}\sum_{i}[\langle k_{i}^{2}\rangle-\text{Var}[k_{i}]-\langle k_{i}\rangle]=\langle S\rangle-\frac{1}{2}\sum_{i}\text{Var}[k_{i}]:$ (45)
@@
the equivalence holding true in case $\text{Var}[k_{i}]=0$, $\forall\>i$ (i.e. either in the microcanonical case or in the canonical, deterministic case). If, instead, the following relationship holds true
@@
In order to solve the system of equations defining the fit2SM, an appropriate vector of initial conditions needs to be chosen. In order to solve the dcGM, we have chosen $z_{0}=1$; in order to solve the fit2SM, we have chosen $\kappa_{i}^{(0)}=\kappa_{i}^{\text{dcGM}}$, $z_{0}=z_{\text{dcGM}}$ and $y_{0}=1$. As a stopping criterium, we have adopted a condition on the infinite norm of the vector of differences between the values of the parameters at subsequent iterations, i.e. $\max\{|\Delta z|,\>|\Delta y|\}<10^{-12}$. The accuracy of our method in estimating the constraints has been evaluated by computing the maximum relative errors defined as $\text{MRE}_{L}=|L^{*}-\langle L\rangle|/L^{*}$ and $\text{MRE}_{S}=|S^{*}-\langle S\rangle|/S^{*}$. Table 1 shows the time employed by our algorithm to converge as well as its accuracy in reproducing the constraints defining it. Overall, our method is fast and accurate: the numerical errors never exceed $O(10^{-1})$ and the time employed to achieve such an accuracy is always less than a minute.
A natural question arises, i.e. does employing the ‘single-iteration’ solution lead to significantly different results? To answer this question, we have compared the ‘single-iteration’ (si) solutions with the self-consistent (sc) ones by calculating the maximum relative error concerning the average BIC values for each time-scale: one finds that $\left|\overline{\text{BIC}}_{\text{sc}}-\overline{\text{BIC}}_{\text{si}}\right|/\overline{\text{BIC}}_{\text{sc}}=0.0016$ at the daily time-scale, $\left|\overline{\text{BIC}}_{\text{sc}}-\overline{\text{BIC}}_{\text{si}}\right|/\overline{\text{BIC}}_{\text{sc}}=0.0011$ at the weekly time-scale, $\left|\overline{\text{BIC}}_{\text{sc}}-\overline{\text{BIC}}_{\text{si}}\right|/\overline{\text{BIC}}_{\text{sc}}=0.0003$ at the monthly time-scale, $\left|\overline{\text{BIC}}_{\text{sc}}-\overline{\text{BIC}}_{\text{si}}\right|/\overline{\text{BIC}}_{\text{sc}}=0.0008$ at the quarterly time-scale, $\left|\overline{\text{BIC}}_{\text{sc}}-\overline{\text{BIC}}_{\text{si}}\right|/\overline{\text{BIC}}_{\text{sc}}=0.0008$ at the yearly time-scale (see also fig. 7).
@@
Let us, now, discuss an alternative way of determining the parameters of the fit2SM. First, let us write the log-likelihood as
@@
and, then, maximise it with respect to $z$ and $y$. Upon doing so, we derive the two, non-linear, coupled equations reading
@@
In case the total number of two-stars were not directly accessible, one could exploit the relationship between $L$ and $S$, replacing $S^{*}$ with the value $S=aL^{b}$. For what concerns eMID, the fitted values read $a_{\text{daily}}\simeq 0.36$, $b_{\text{daily}}\simeq 1.59$; $a_{\text{weekly}}\simeq 0.36$, $b_{\text{weekly}}\simeq 1.61$; $a_{\text{monthly}}\simeq 0.47$, $b_{\text{monthly}}\simeq 1.59$; $a_{\text{quarterly}}\simeq 0.67$, $b_{\text{quarterly}}=1.56$; $a_{\text{yearly}}\simeq 0.69$, $b_{\text{yearly}}=1.56$.
Figure 8: Top panel: distribution of the spectral radius induced by $M=10^{3}$ configurations sampled from the Chung-Lu model defined by the topology of the eMID snapshot corresponding to the day 20102010-0707-1919. Bottom panels: estimations of $\lambda_{1}$, obtained from each of the $M$, sampled configurations, scattered versus the corresponding empirical values, for the UBCM (red), the dcGM (yellow) and the fit2SM (brown). Vertical bars indicate the standard deviation accompanying the estimation carried out on the specific configuration. While the dcGM generally underestimates the spectral radius of the generative model and the UBCM overestimates it, the fit2SM captures it quite accurately.
@@
Figure 9: Comparison between the ensemble averages of the total number of isolated nodes with their analytical counterparts for the weekly, monthly, quarterly, yearly aggregation levels in 1999 (left) and 2012 (right): overall, the estimations provided by the UBCM (red), the dcGM (yellow) and the fit2SM (brown) are very accurate for each time-scale. The quarter, month, week and day shown for each year correspond to those reported in Table 1.
Figure 10: Analysis of eMID at (from top-left to bottom-right) the daily, weekly, monthly, quarterly and yearly time-scales. According to the expected number of isolated nodes, the fit2SM outperforms the dcGM at the daily and weekly time-scales and compete with the dcGM and the UBCM at the monthly and quarterly time-scales. Trends have been smoothed via a rolling average over the points [t10,t+10][t-10,t+10] for the daily aggregation, [t7,t+7][t-7,t+7] for the weekly aggregation, [t5,t+5][t-5,t+5] for the monthly aggregation, [t3,t+3][t-3,t+3] for the quarterly aggregation.
Let us, now, compare
@@
across our temporal snapshots. Figure 9 compares the ensemble averages of the total number of isolated nodes with their analytical counterparts for the weekly, monthly, quarterly, yearly aggregation levels in 1999 and 2012: overall, the estimations provided above are very accurate at each time-scale. Figure 10, instead, compares the performances of the UBCM, the dcGM and the fit2SM in ‘producing’ configurations being characterised by a certain number of isolated nodes: as we have already observed, the sparser the configuration, the better the performance of the fit2SM - in these cases, the fit2SM outperforms the dcGM and compete with the UBCM on at least a portion of the dataset, by allowing a smaller number of isolated nodes to appear on sampled configurations.

### v4

@@ VI.3 Reproducing a network structure
A compact indicator of a model’s performance in reconstructing a network structure is provided by the Bayesian Information Criterion (BIC), reading
@@
where V=N(N1)/2V=N(N-1)/2 accounts for the system size, the number of parameters amounts at $c=N$ for the UBCM, $c=1$ for the dcGM, and $c=2$ for the fit2SM, and the log-likelihood reads
@@
$\delta\text{BIC}_{\text{model}}=\frac{\text{BIC}_{\text{model}}-\text{BIC}_{\text{min}}}{\text{BIC}_{\text{model}}},$ (27)
where $\text{BIC}_{\text{min}}=\min_{\text{model}}\{\text{BIC}_{\text{model}}\}$; while a null value of such an index identifies the best-performing model on the considered snapshot, the larger its value, the worse the performance of the corresponding model. As fig. 4 shows, our canonical, non-linear model
  • steadily outperforms both the dcGM and the (much) more demanding UBCM at the daily and weekly time-scales, i.e., on sparser configurations (in fact, $\text{BIC}_{\text{fit2SM}}<\text{BIC}_{\text{dcGM}}$ on $90.76\%$ of daily snapshots and $97.63\%$ of weekly snapshots; $\text{BIC}_{\text{fit2SM}}<\text{BIC}_{\text{UBCM}}$ on $99.97\%$ of daily snapshots and $100.00\%$ of weekly snapshots);
@@
Overall, these findings suggest the presence of a dependency between the performance of the fit2SM and the network link density, i.e., the sparser the network, the better its performance. To test such an hypothesis, let us scatter the index of relative performance
$\Delta\text{BIC}_{\text{model}}=\frac{\text{BIC}_{\text{model}}-\text{BIC}_{\text{fit2SM}}}{\text{BIC}_{\text{model}}}$ (28)
@@
$\Delta S_{\text{model}}=\frac{\langle S\rangle_{\text{model}}-\langle S\rangle_{\text{fit2SM}}}{\langle S\rangle_{\text{model}}}=\frac{\langle S\rangle_{\text{model}}-S^{*}}{\langle S\rangle_{\text{model}}}$ (29)
for both the UBCM and the dcGM. As fig. 5 reveals, whenever $\Delta S_{\text{model}}$ is large - typically the case for sparser configurations - model selection favours the fit2SM.
@@ VII Discussion
After reviewing a number of negative results about the possibility of defining an ERG constraining both the degree sequence and the total number of two-stars, we propose a minimal, canonical model capable of reproducing the variance of a degree distribution while accounting for the nodes heterogeneity: in order to achieve such a goal, a fitness-induced variant of the 2SM, named fit2SM (i.e., a ‘softened’ non-linear ERG whose definition is inspired to its fitness-based, linear counterpart), has been employed.
Besides achieving a large accuracy in reproducing a network structure at different levels, the fit2SM seems also to mitigate a problem affecting the dcGM, i.e. that of returning sampled configurations with too many isolated nodes: in fact, $\langle N_{\text{fit2SM}}^{0}\rangle<\langle N_{\text{dcGM}}^{0}\rangle$ on $97.63\%$ of daily snapshots, $98.61\%$ of weekly snapshots, and $67.88\%$ of monthly snapshots; interestingly, $\langle N_{\text{fit2SM}}^{0}\rangle<\langle N_{\text{UBCM}}^{0}\rangle$ on $58.79\%$ of monthly snapshots and $38.10\%$ of quarterly snapshots but $\langle N_{\text{fit2SM}}^{0}\rangle<\langle N_{\text{UBCM}}^{0}\rangle$ on $66.67\%$ of monthly snapshots and $55.56\%$ of quarterly snapshots since 2009 (see also Appendix F).
@@
Future research will explore non-linear models enforcing more complex patterns (e.g., the Strauss one, constraining the total number of triangles).
@@ X Acknowledgments
MM, DG and TS acknowledge support from the project ‘SoBigData.it - Strengthening the Italian RI for Social Mining and Big Data Analytics’ - IR0000013 - CUP B53C22001760006, financed by European Union - Next Generation EU - National Recovery and Resilience Plan (Piano Nazionale di Ripresa e Resilienza, PNRR) - M4C2 I.3.1; DG and TS acknowledge support from the projects ‘Reconstruction, Resilience and Recovery of Socio-Economic Networks’ RECON-NET EP_FAIR_005 - PE0000013 ‘FAIR’, Funded by the European Union Next Generation EU, PNRR Mission 4 Component 2, Investment 3.1; ‘RE-Net: Reconstructing economic networks: from physics to machine learning and back’ - 2022MTBB22, Funded by the European Union Next Generation EU, PNRR Mission 4 Component 2 Investment 1.1, CUP: D53D23002330006; ‘C2T - From Crises to Theory: towards a science of resilience and recovery for economic and financial systems’ - P2022E93B8, Funded by the European Union Next Generation EU, PNRR Mission 4 Component 2 Investment 1.1, CUP: D53D23019330001. This work is also supported by the European Union - NextGenerationEU and funded by the Italian Ministry of University and Research (MUR) - National Recovery and Resilience Plan (Piano Nazionale di Ripresa e Resilienza, PNRR), project ‘A Multiscale integrated approach to the study of the nervous system in health and disease’ (MNESYS, PE0000006, DN. 1553, 11/10/2022). We thank Giulio Cimini for fruitful and insightful discussions.
@@ XII Competing Interests
We employ transaction-level records from the Electronic Market for Interbank Deposits (eMID), a screen-based market for unsecured deposits [8, 37, 38], restricting the analysis to the segment that represents the vast majority of the activity on such a market, i.e., the overnight one. Our sample spans trading days from January $1999$ to September $2012$: each trading day $d$ defines a weighted, directed matrix $\mathbf{V}(d)$ whose entry $v_{ij}(d)$ equals the total notional amount lent by bank $i$ to bank $j$ on day $d$; naturally, $v_{ii}(d)=0$.
Throughout the paper, the labels ‘daily’, ‘weekly’, ‘monthly’, ‘quarterly’ and ‘yearly’ refer to calendar aggregations of daily records: more specifically, ‘weekly’ refers to ISO calendar weeks (Monday to Sunday), ‘monthly’ to calendar months, ‘quarterly’ to standard quarters (Q1: Jan-Mar; Q2: Apr-Jun; Q3: Jul-Sep; Q4: Oct-Dec) and ‘yearly’ to calendar years. Importantly, trading on eMID only occurs on business days, so that weekends and bank holidays do not contribute to observations: consequently, the number of trading days within a given calendar window is not constant across windows of the same kind; in practice, we aggregate over the set of trading days that are present in the raw records during a specific window. Let $\Delta_{t}$ denote the set of trading days belonging to the calendar window indexed by $t$ (week, month, quarter, or year). We, thus, aggregate weights according to
@@
as a last observation, let us stress that, for each window, we restrict the set of nodes to the banks that are active within that window, i.e., that are involved in at least one transaction during $\Delta_{t}$: the number of nodes, thus, becomes a time-dependent quantity.
@@
with $a_{ii}^{\Delta}(t)=0$. Naturally, we compute the node strengths from the underlying, symmetrized weights as
@@
To ease the mathematical manipulations, let us focus on binary, undirected networks. Linear ERGs are described by Hamiltonians reading
@@
hence inducing models that can be factorized as
@@
hence inducing models that can be factorized as well as
@@
where $\mathbf{P}=\langle\mathbf{A}\rangle$. The mean-field approximation consists precisely in replacing $\mathbf{A}$ with $\mathbf{P}$ in $f_{ij}$, in eq. 36.
@@
$\sum_{i}\binom{\langle k_{i}\rangle}{2}=\frac{1}{2}\sum_{i}[\langle k_{i}\rangle^{2}-\langle k_{i}\rangle]=\frac{1}{2}\sum_{i}[\langle k_{i}^{2}\rangle-\text{Var}[k_{i}]-\langle k_{i}\rangle]=\langle S\rangle-\frac{1}{2}\sum_{i}\text{Var}[k_{i}].$ (45)
@@
the equivalence holding true in case $\text{Var}[k_{i}]=0$, $\forall\>i$ (i.e., either in the microcanonical case or in the canonical, deterministic case). If, instead, the following relationship holds true:
@@
In order to solve the system of equations defining the fit2SM, an appropriate vector of initial conditions needs to be chosen. In order to solve the dcGM, we have chosen $z_{0}=1$; in order to solve the fit2SM, we have chosen $\kappa_{i}^{(0)}=\kappa_{i}^{\text{dcGM}}$, $z_{0}=z_{\text{dcGM}}$, and $y_{0}=1$. As a stopping criterium, we have adopted a condition on the infinite norm of the vector of differences between the values of the parameters at subsequent iterations, i.e., $\max\{|\Delta z|,\>|\Delta y|\}<10^{-12}$. The accuracy of our method in estimating the constraints has been evaluated by computing the maximum relative errors defined as $\text{MRE}_{L}=|L^{*}-\langle L\rangle|/L^{*}$ and $\text{MRE}_{S}=|S^{*}-\langle S\rangle|/S^{*}$. Table 1 shows the time employed by our algorithm to converge as well as its accuracy in reproducing the constraints defining it. Overall, our method is fast and accurate: the numerical errors never exceed $O(10^{-1})$ and the time employed to achieve such an accuracy is always less than a minute.
A natural question arises, i.e., does employing the ‘single-iteration’ solution lead to significantly different results? To answer this question, we have compared the ‘single-iteration’ (si) solutions with the self-consistent (sc) ones by calculating the maximum relative error concerning the average BIC values for each time-scale: one finds that $\left|\overline{\text{BIC}}_{\text{sc}}-\overline{\text{BIC}}_{\text{si}}\right|/\overline{\text{BIC}}_{\text{sc}}=0.0016$ at the daily time-scale, $\left|\overline{\text{BIC}}_{\text{sc}}-\overline{\text{BIC}}_{\text{si}}\right|/\overline{\text{BIC}}_{\text{sc}}=0.0011$ at the weekly time-scale, $\left|\overline{\text{BIC}}_{\text{sc}}-\overline{\text{BIC}}_{\text{si}}\right|/\overline{\text{BIC}}_{\text{sc}}=0.0003$ at the monthly time-scale, $\left|\overline{\text{BIC}}_{\text{sc}}-\overline{\text{BIC}}_{\text{si}}\right|/\overline{\text{BIC}}_{\text{sc}}=0.0008$ at the quarterly time-scale, and $\left|\overline{\text{BIC}}_{\text{sc}}-\overline{\text{BIC}}_{\text{si}}\right|/\overline{\text{BIC}}_{\text{sc}}=0.0008$ at the yearly time-scale (see also fig. 7).
@@
Let us now discuss an alternative way of determining the parameters of the fit2SM. First, let us write the log-likelihood as
@@
and, then, maximize it with respect to $z$ and $y$. Upon doing so, we derive the two, non-linear, coupled equations reading
@@
In case the total number of two-stars were not directly accessible, one could exploit the relationship between $L$ and $S$, replacing $S^{*}$ with the value $S=aL^{b}$. For what concerns eMID, the fitted values read $a_{\text{daily}}\simeq 0.36$, $b_{\text{daily}}\simeq 1.59$; $a_{\text{weekly}}\simeq 0.36$, $b_{\text{weekly}}\simeq 1.61$; $a_{\text{monthly}}\simeq 0.47$, $b_{\text{monthly}}\simeq 1.59$; $a_{\text{quarterly}}\simeq 0.67$, $b_{\text{quarterly}}=1.56$; and $a_{\text{yearly}}\simeq 0.69$, $b_{\text{yearly}}=1.56$.
Figure 8: Top panel: distribution of the spectral radius induced by $M=10^{3}$ configurations sampled from the Chung-Lu model defined by the topology of the eMID snapshot corresponding to the day 20102010-0707-1919. Bottom panels: estimations of $\lambda_{1}$, obtained from each of the $M$, sampled configurations, scattered versus the corresponding empirical values, for the UBCM (red), the dcGM (yellow), and the fit2SM (brown). Vertical bars indicate the standard deviation accompanying the estimation carried out on the specific configuration. While the dcGM generally underestimates the spectral radius of the generative model and the UBCM overestimates it, the fit2SM captures it quite accurately.
@@
Figure 9: Comparison between the ensemble averages of the total number of isolated nodes with their analytical counterparts for the weekly, monthly, quarterly, and yearly aggregation levels in 1999 (left) and 2012 (right): overall, the estimations provided by the UBCM (red), the dcGM (yellow), and the fit2SM (brown) are very accurate for each time-scale. The quarter, month, week, and day shown for each year correspond to those reported in Table 1.
Figure 10: Analysis of eMID at (from top-left to bottom-right) the daily, weekly, monthly, quarterly, and yearly time-scales. According to the expected number of isolated nodes, the fit2SM outperforms the dcGM at the daily and weekly time-scales and compete with the dcGM and the UBCM at the monthly and quarterly time-scales. Trends have been smoothed via a rolling average over the points [t10,t+10][t-10,t+10] for the daily aggregation, [t7,t+7][t-7,t+7] for the weekly aggregation, [t5,t+5][t-5,t+5] for the monthly aggregation, [t3,t+3][t-3,t+3] for the quarterly aggregation.
Let us now compare
@@
across our temporal snapshots. Figure 9 compares the ensemble averages of the total number of isolated nodes with their analytical counterparts for the weekly, monthly, quarterly, and yearly aggregation levels in 1999 and 2012: overall, the estimations provided above are very accurate at each time-scale. Figure 10, instead, compares the performances of the UBCM, the dcGM and the fit2SM in ‘producing’ configurations being characterized by a certain number of isolated nodes: as we have already observed, the sparser the configuration, the better the performance of the fit2SM - in these cases, the fit2SM outperforms the dcGM and compete with the UBCM on at least a portion of the dataset, by allowing a smaller number of isolated nodes to appear on sampled configurations.